A catering service offes 9 appetizers, 12 main courses, and 7 desserts. A banquet chairperson is to select 6 appetizers, 5 main courses, and 2 desserts for a banquet. In how many ways can this be done?

Answer by Boreal(15235)

This is 9C6*12C5*7C2=84*792*21=1,397,088